# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

### v1

@@ 1. Introduction
due to Huxley [Hux96]  [^1] Recently, Xiaochun Li and Xuerui Yang have posted an article (see https://arxiv.org/abs/2308.14859) that claims to improve this bound of Huxley., while the conjectured bound $O(r^{1/2+\varepsilon})$ remains hopelessly out of reach of the current technology. If we replace the circle by a hyperbola $xy=N$, with $N\longrightarrow\infty$, the problem of counting lattice points with positive coordinates under the hyperbola is known as Dirichlet’s divisor problem and the situation here is completely analogous to the Gauss circle problem (see [IK04, Chap. 1]). An interesting lattice point counting problem concerns counting elements $\gamma\in SL(2,{\mathbb{Z}})$; i.e., integer matrices $\gamma=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ with determinant $ad-bc=1$, such that $||\gamma||\leq x$, where we take the norm to be the standard Euclidean norm on ${{\mathbb{R}}}^{4}$ after identifying the matrix $\gamma$ with the point $(a,b,c,d)$; i.e.,
$||\gamma||^{2}=a^{2}+b^{2}+c^{2}+d^{2}.$
@@
then $d(\gamma.i,i)=a^{2}+b^{2}+c^{2}+d^{2}$, where $\gamma.z$ denotes the usual Möbius action $\gamma.z=(az+b)(cz+d)^{-1}.$ Thus the above problem is equivalent to studying the asymptotic behaviour of the size of the set $\{\gamma\in\Gamma:d(\gamma.i,i)\leq x\}$, where $\Gamma$ is the group $SL(2,{\mathbb{Z}})$ and one may also consider a more general problem by taking $\Gamma$ to be any discrete subgroups of $SL(2,{\mathbb{R}})$ with finite co-volume. Unlike in the Euclidean case, even finding the asymptotic for lattice point counting problems is not easy and the elementary counting arguments no longer work due to the negative curvature of the underlying space. Lattice point counting problems are intimately related to the spectrum of the Laplace operator on appropriate manifolds. For example, the Gauss circle problem is nothing but the problem of improving the bound on the error term in the Weyl law for the flat torus ${{\mathbb{Z}}}^{2}\backslash{{\mathbb{R}}}^{2}$. The above hyperbolic lattice point counting problem is related to the spectrum of the hyperbolic Laplacian acting on $\ΓL^{2}(\Gamma\backslash\mathbb{H})$ (as a densely defined operator) and after initial advances made by Delsarte [Del42] and Huber, who solved the problem for cocompact groups (see [Hub56, Hub61]), the original problem for $\Gamma=SL(2,{\mathbb{Z}})$ could be tackled once Selberg established the complete spectral resolution of the Laplacian for groups with parabolic elements. Selberg’s result (see, e.g., [IK04, Corollary 15.12]) says
$\sum_{\begin{subarray}{c}\gamma\in SL(2,{\mathbb{Z}})\\ ||\gamma||\leq x\end{subarray}}1=6x^{2}+O(x^{4/3}).$ (1.1)
For an account of this beautiful and deep circle of ideas, see, e.g., [IK04, Chap. 15], [Iwa02], [LP82], [Pat75] and the original articles by Selberg, viz. [Sel56], [Sel63].
This result has been vastly generalized by Duke, Rudnick and Sarnak [DRS93] to varieties $V$ with an algebraic group action for which $V(\mathbb{R})$ is affine and symmetric by applying advanced results from harmonic analysis. A special case of their result says
$\sum_{\begin{subarray}{c}\gamma\in M(n,{\mathbb{Z}})\\ det(\gamma)=r\\ \\ ||\gamma||\leq x\end{subarray}}1=c_{n,r}x^{n^{2}-n}+O_{\varepsilon}(x^{n(n-1)-1/(n+1)+\varepsilon}),$ (1.2)
@@
$\mathpzc{S}_{r}(X)=\frac{16}{\zeta(2)}\frac{\sigma(|r|)}{|r|}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}),$ (1.3)
as long as $r=O\left(X^{{1/3}}\right)$. Here $\theta$ is any admissible exponent for the Generalized Ramanujan Conjecture, $\varepsilon$ is an arbitrarily small positive real number, and $\sigma(n)$ denotes the sum of the divisors of a positive integer $n$.
By the work of Kim and Sarnak [Kim03, Appendix 2] it is known that $\theta\leq\frac{7}{64}$.
@@
The reason we need to take $r=O(X^{1/3})$ is explained in Remark 16.1. However, with more work, it seems possible to extend the range of $r$. Also, at the cost of a worse exponent of $X$, we can make the dependence of $r$ on the error term disappear. See Remark 7.6 for a precise statement.
As mentioned earlier, we can put our problem in the general context of counting integral points on affine hypersurfaces defined over ${\mathbb{Z}}$ and more specifically to counting vectors $\mathbf{a}\in{{\mathbb{Z}}}^{n}$ to equations of the form
$Q(\mathbf{x})=r$ (1.4)
in an expanding box $[-X,X]^{n^{2}}$, where $Q$ is a quadratic form in $n$ variables over ${\mathbb{Z}}$ and $r$ is a fixed integer. This is a classical problem with a rich history and the circle method, developed and refined by Ramanujan, Hardy, Littlewood, Kloosterman and many others, has been employed here with great success. See the article [HB96] by Heath-Brown for an introduction with modern perspectives to this problem. Usually one counts integral solutions to (1.4) by attaching weights $w_{i}(a_{i}/X)$, where $w_{i}$’s are smooth functions with compact support and it is possible to get fairly strong bound on the error term relative to the main term in many cases (see, e.g., [HB96, Theorems 4-8]). If we count with sharp cuts, however, then the situation is very different and in general we can only get a saving of an arbitrary power of $\log X$ (see [Nie10]). It is interesting, therefore, to address this problem for forms of special shapes in a small number of variables and the determinant form $Q(x,y,z,w)=xy-zw$ is particularly suitable because of the structural advantage it affords. In fact, the only two results we could find in the literature that deal with unweighted counting are about determinant forms in two or more variables. Gorodnik and Nevo (see [GN10, Cor. 2.3]) studied the problem of counting matrices in $SL(n,{\mathbb{Z}})$ inside a ball with respect to a general norm using ergodic method and in the case $n=2$, the saving in the exponent in the error term is $1/16$. Roettger [Roe05] obtained an asymptotic formula for the number of unimodular matrices over the ring of integers in a number field where the elements have height up to some parameter $t$. When specialized to our case $K={\mathbb{Q}}$, the error term gives a saving of $1/15$ in the exponent. In comparison, we save $1/2$ in the exponent of the error term in Theorem 1.1.
If we count with smooth weights then we can save even more and can also take $r$ to be much larger as the following result shows.
Theorem 1.3.
Let
$S_{V}(X,r)=\mathop{\sum_{a}\sum_{b}\sum_{c}\sum_{d}}_{\begin{subarray}{c}ad-bc=r\end{subarray}}V\left(\frac{a}{X}\right)V\left(\frac{b}{X}\right)V\left(\frac{c}{X}\right)V\left(\frac{d}{X}\right),$
where $V$ is a fixed smooth function with compact support in $[1,2]$. Then for $0<r\ll X^{2}$, we have
$S_{V}(X,r)=M_{V}(X,r)+O_{\varepsilon}\left(r^{1/4}X^{1+\varepsilon}\right),$
@@
$M_{V}(X,r)=\sum_{l|r}\sum_{k>0}\frac{\mu(k)}{k}\iiint\frac{1}{z}V\left(\frac{x}{X}\right)V\left(\frac{lky}{X}\right)V\left(\frac{lkz}{X}\right)V\left(\frac{r+lkxy}{zlkX}\right)\textnormal{d}x\textnormal{d}y\textnormal{d}z.$
It is not hard to see that $M_{V}(X,r)\asymp X^{2}$ if $r$ is not very large; e.g., if $0<|r|\leq 2X^{2}$. Moreover, by the mean value theorem of Calculus, we see that
$S_{V}(X,r)=KX^{2}+O\left(\sigma(r)\right)+O\left(r^{1/4}X^{1+\varepsilon}\right),$
where
$K=\frac{1}{\zeta(2)}\frac{\sigma(r)}{r}\iiint\frac{1}{t}V\left(u\right)V\left(v\right)V\left(t\right)V\left(\frac{uv}{t}\right)\textnormal{d}u\,\textnormal{d}v\,\textnormal{d}t.$
Although we use automorphic methods to prove this result, a result of comparable strength can also be obtained by employing the method of [HB96]. In fact, if we take $r$ to be of size $X^{2}$ then Theorem 1.3 gives an error term of the same strength as the one in [HB96, Theorem 4]. However, it is possible to improve the exponent of $r$ to $\theta$ with more work. It is of interest to consider more general quadratic forms and to investigate whether some of the results of [HB96] (see also [Get18]) can be refined in special cases by introducing the spectral methods of automorphic forms into the analysis after the application of the $\delta$-symbol method. In this context, see also the discussion after [HB96, Cor. 1]. We have included the above result to illustrate the impact of counting with smooth weights and have given only a brief sketch of the proof in §1.2 since this proof is much easier compared to the proof of Theorem 1.1. An interesting variation of the general counting problem is to attach arithmetic weights such the von Mangold function or coefficients of cusp forms to the variables of summation (see [Kum20]) and this is the topic of an upcoming work by the second author along with P. Sharma and S.K. Singh.
1.1. Extensions and applications of the results
1.1.1. Determinant equations in Analytic Number Theory
Although we are mainly motivated by the intrinsic interest in the problem, counting solutions to the determinant equations $ad-bc=r$, $r\neq 0$ with certain weights $\alpha(a),\alpha(b)$ etc. come up naturally in the subject, for example, in the study of amplified moments of $L$-functions (see, e.g., [DFI97], [DFI02, §16], [BC18]). In [BC18] and [DFI97], sums of the form
@@ 1.1.1. Determinant equations in Analytic Number Theory
were studied and asymptotic formulae for such sums were obtained. Here, $\mathcal{N}_{i}=[N_{i}/2,N_{i}]$, $\alpha,\beta$ are arbitrary sequences, and $f,g$ are smooth functions satisfying suitable decay conditions on their derivatives. It is possible to allow arbitrary weights in some of the variables at the cost of an worse error term. For example, the following is a consequence of an intermediate step in the proof of Theorem 1.1.
@@
Unlike (1.5), here only one variable is weighted by an arbitrary sequence.
If one chooses weights of arithmetic nature then further possibilities for improvements may open up. For example, if we take $\alpha(a)=\mathbf{1}_{a\equiv 0(\text{mod }q)}$, the indicator function of the integral multiples of $q$, where $q>1$ is a fixed integer and keep the other weights the constant function $1$ (or smooth weights $W(a/X),W(b/X)$ etc.), then also our method applies and this leads to the study of the spectral theory for the congruence group $\Gamma_{0}(q)$.
1.1.2. Extensions to other quadratic forms in four variables
The method of our proof does not distinguish between the determinant equation $xy-zw=r$ and the equation $pxy-qzw=r$ where $p,q$ are fixed integers such that $(p,q)|r$. One can easily obtain analogous results for the latter equation of the same strength as long as $p,q=O(X^{\varepsilon})$. In fact, by a simple change of variables, we can derive an asymptotic formula for the number of integer solutions to an equation of the form $f_{1}(x,y)-f_{2}(z,w)=r$, where $f_{1}$ and $f_{2}$ both factor over ${\mathbb{Z}}$ as products of two linear forms and the coefficients of both forms are assumed to be $O(X^{\varepsilon})$. However, the formula we obtain in this way counts the number of solutions in a new domain that depends on the coefficients of $f_{1}$ and $f_{2}$. To clarify the situation, let us give an illustrative example. Let us take $f_{1}(x,y)=f_{2}(x,y)=x^{2}-y^{2}$. Then we obtain the surface $x^{2}-y^{2}-z^{2}+w^{2}=r$. Now, by the change of variables $(x,y,z,w)\rightarrow(x^{\prime},y^{\prime},z^{\prime},w^{\prime})$, where $x+y=x^{\prime},x-y=y^{\prime},z+w=z^{\prime},z-w=z^{\prime}$, we obtain the surface $x^{\prime}y^{\prime}-z^{\prime}w^{\prime}=r$. By Theorem 1.1, we have a counting formula for the number of integer solutions to $x^{\prime}y^{\prime}-z^{\prime}w^{\prime}=r$, where $(x^{\prime},y^{\prime},z^{\prime},w^{\prime})\in[-X,X]^{4}$. Now the inverse change of variables transform this box to a four-dimensional tetrahedron $T(X)$ which can be described as the convex hull of the points
$(X,0,0,0),(-X,0,0,0),(0,X,0,0),(0,-X,0,0),\cdots,(0,0,0,X),(0,0,0,-X).$
Thus we obtain the following asymptotic formula:
Corollary 1.5.
We have,
$\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}a^{2}-b^{2}-c^{2}+d^{2}=r\\ a,b,c,d\in T(X)\end{subarray}}1=\frac{16}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}).$
1.1.3. Extensions to quadratic forms in three variables
If we consider quadratic forms of the shape $Q(x,y,z)=px^{2}-qyz$, then also our methods can be adopted to study integral solutions to $Q(x,y,z)=r$. Here, one encounters a sum of Salié sums instead of Kloosterman sums and Proskurin’s formula generalizing the Kuznetsov formula (see [Pro03]) is applicable here.
1.1.4. Relation to the shifted convolution sum problem for the restricted divisor function
The shifted convolution sum refers to the problem of the estimation or obtaining asymptotic formula for sums of the type $\sum_{n\leq x}\lambda(n)\lambda(n+h)$ or the corresponding smoothed sum, where $\lambda$ is an interesting arithmetical function, e.g., the divisor function, coefficients of a cusp form etc., and the shift $h$ is a fixed non-zero integer. Lattice point counting problems can sometimes be interpreted as the shifted convolution problem for coefficients of appropriate automorphic forms. For example, the formula (1.1) is equivalent to the formula
@@ 1.1.4. Relation to the shifted convolution sum problem for the restricted divisor function
Compared to the classical additive divisor problem, estimating the error term for this sum is harder and the reason is clear. Writing $n=bc$ and $n+r=ad$, here all the variables are individually restricted by sharp edges, whereas in the sum $\sum_{n\leq X}\tau(n)\tau(n+h)$, the four variables are subject to only one such restriction ($bc\leq X$) and it is possible to treat two of the three independent variables as “smooth” variables.
1.2. Method of the proofs of Theorem 1.1 and Theorem 1.3
The theorems above are proved by the so called “Kloostermania” techniques, i.e., we reduce the problems to that of estimation of sums of Kloosterman sums which is then carried out by employing the spectral theory of automorphic forms. Let us first consider the proof of Theorem 1.3 which serves as a warm-up for the proof of Theorem 1.1. The first step is to eliminate one of the four variables, say $d$, and to interpret the determinant equation as a congruence modulo one of the other variables, say, $a$. This is a natural and well-known idea. Then we apply Poisson summation (see Lemma 2.1) to one of the remaining two variables, say $b$. As usual, the 00-th frequency in the dual sum gives the main term and after one more application of Poisson summation in the variable $c$ to the remaining sum yields a sum of Ramanujan sums (coming from the $zero$-th term in new dual variable) that can be easily estimated, plus the following double sum:

### v2

@@ 1. Introduction
due to Huxley [Hux96]  [^1] Recently, Xiaochun Li and Xuerui Yang have announced (see https://arxiv.org/abs/2308.14859) an improvement over the bound of Huxley. and the conjectured bound $O(r^{1/2+\varepsilon})$ remains hopelessly out of reach of the current technology. If we replace the circle by a hyperbola $xy=N$, with $N\longrightarrow\infty$, the problem of counting lattice points with positive coordinates under the hyperbola is known as Dirichlet’s divisor problem and the situation here is completely analogous to the Gauss circle problem (see [IK04, Chap. 1]). An interesting lattice point counting problem concerns counting elements $\gamma\in SL(2,{\mathbb{Z}})$; i.e., integer matrices $\gamma=\begin{pmatrix}a&b\\ c&d\\ \end{pmatrix}$ with determinant adbc=1ad-bc=1, such that $\gamma||\leq x$, where we take the norm to be the standard Euclidean norm on ${{\mathbb{R}}}^{4}$ after identifying the matrix $\gamma$ with the point $(a,b,c,d)$; i.e.,
$\gamma||^{2}=a^{2}+b^{2}+c^{2}+d^{2}.$
@@
then $d(\gamma.i,i)=a^{2}+b^{2}+c^{2}+d^{2}$, where $\gamma.z$ denotes the usual Möbius action $\gamma.z=(az+b)(cz+d)^{-1}.$ Thus the above problem is equivalent to studying the asymptotic behaviour of the size of the set $\{\gamma\in\Gamma:d(\gamma.i,i)\leq x\}$, where $\Gamma$ is the group $SL(2,{\mathbb{Z}})$ and one may also consider a more general problem by taking $\Gamma$ to be any discrete subgroups of $SL(2,{\mathbb{R}})$ with finite co-volume. Unlike in the Euclidean case, even finding the asymptotic for lattice point counting problems is not easy and the elementary counting arguments no longer work due to the negative curvature of the underlying space. Lattice point counting problems are intimately related to the spectrum of the Laplace operator on appropriate manifolds. For example, the Gauss circle problem is nothing but the problem of improving the bound on the error term in the Weyl law for the flat torus ${{\mathbb{Z}}}^{2}\backslash{{\mathbb{R}}}^{2}$. The above hyperbolic lattice point counting problem is related to the spectrum of the hyperbolic Laplacian acting on $\)L^{2}(\Gamma\backslash\mathbb{H}$) (as a densely defined operator) and after initial advances made by Delsarte [Del42] and Huber, who solved the problem for cocompact groups (see [Hub56, Hub61]), the original problem for $\Gamma=SL(2,{\mathbb{Z}})$ could be tackled once Selberg established the complete spectral resolution of the Laplacian for groups with parabolic elements. Selberg’s result (see, e.g., [IK04, Corollary 15.12]) says
@@
For an account of this beautiful and deep circle of ideas, see, e.g., [IK04, Chap. 15], [Iwa02], [LP82], [Pat76] and the original articles by Selberg, viz. [Sel56], [Sel63]. For an interesting twisted version of Selberg’s result, see [Goo83]. Selberg’s work has been vastly generalized by Duke, Rudnick and Sarnak [DRS93] to varieties $V$ with an algebraic group action for which $V(\mathbb{R})$ is affine and symmetric by applying advanced results from harmonic analysis. A special case of their result says
[…]
where $M(n,{\mathbb{Z}})$ is the set of $n\times n$ integer matrices, $\|\cdot\|$ denotes the Euclidean norm $\sqrt{tr(\gamma{\gamma}^{t})}$, $r\neq 0$ is a fixed integer, (i.e., $r=O(1)$), $c_{n,r}>0$ is an explicit constant and $\varepsilon>0$ is arbitrary. The saving $\delta_{n}=1/(n+1)$ in the exponent of the error term in (1.2) was later improved by Gorodnik, Nevo and Yehoshua [GNY17] to, roughly, $2\delta_{n}$ and very recently it has been improved further by Blomer and Lutsko [BL24] who have shown that the saving is at least $1+1/\sqrt{2}+O(1/n)$ (for $r=1$ and $n\geq 3$). In the case $n=2$, however, the bound $O(X^{4/3})$ for the error term due to Selberg remains unbeaten to this day.
@@
$\mathcal{S}_{r}(X)=\frac{16}{\zeta(2)}\frac{\sigma(|r|)}{|r|}X^{2}+O_{\varepsilon}(|r|^{\theta}X^{3/2+\varepsilon}),$ (1.3)
as long as $r=O\left(X^{{1/3}}\right)$. Here $\theta$ is an admissible exponent towards the Ramanujan Conjecture.
Conjecturally $\theta$ can be taken to be zero and by the work of Kim and Sarnak [Kim03, Appendix 2] it is known that $\theta\leq\frac{7}{64}$. Here $\varepsilon$ is any positive real number that can be arbitrarily small and $\sigma(n)$ denotes the sum of all divisors of a positive integer $n$.
@@
The question we address here can be put in the general context of counting integral points on surfaces of the type $Q(x1,x_{2},\cdots,x_{n})=r$ where $Q$ is a quadratic form. See [HB96] for an introduction to this topic. However, counting with smooth weights seems more natural from that point of view and we have considered the “smooth” analogue of this problem in our companion article [GG25]. The interest in our question from this angle has been discussed in some detail in the introduction to [GG25]. As one should expect, we get a much stronger error term in this situation. Indeed, if $V$ is a fixed smooth function with compact support, say, in $[1,2]$ and if $r$ is a fixed non-zero integer, then we show
$\mathop{\sum_{a}\sum_{b}\sum_{c}\sum_{d}}_{\begin{subarray}{c}ad-bc=r\end{subarray}}V\left(\frac{a}{X}\right)V\left(\frac{b}{X}\right)V\left(\frac{c}{X}\right)V\left(\frac{d}{X}\right)=M_{V}(X,r)+O_{\varepsilon}\left(|r|^{\theta}X^{1+\varepsilon}\right),$
@@
$M_{V}(X,r)=\sum_{l|r_{1}}\sum_{k>0}\frac{\mu(k)}{k}\iiint\frac{1}{z}V\left(\frac{x}{X}\right)V\left(\frac{lky}{\ X}\right)V\left(\frac{lkz}{X}\right)V\left(\frac{r_{1}+lkxy}{zlkX}\right)\textnormal{d}x\textnormal{d}y\textnormal{d}z,$
and for $0\neq r\in(-3X^{2},X^{2})$, we may write $M_{V}(X,r)=c_{V}\frac{\sigma(|r|)}{|r|}X^{2}+O(\sigma(|r|)$, where $c_{V}>0$ is a constant depending only on $V$.
The proof of this result, though non-trivial, is considerably simpler than the proof of Theorem 1.3. See §1.2 below or [GG25] for further details.
Remark 1.2.
The reason we need to take $r=O(X^{1/3})$ is explained in Remark 15.1. This restriction can possibly be relaxed with more work but we have not tried to do so in this paper. Also, at the cost of a worse exponent of $X$, we can enlarge the admissible range of $r$. In particular, we have
${\mathcal{S}_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(\max\left(X^{5/3},|r|\right)X^{\varepsilon}\right).$
See Remark 3.2 for a proof. Muhammad Afifurrahman [Afi24] has also independently obtained the same result recently by a totally different method. We employ an elaborate stationary phase analysis and the Weil bound for Kloosterman sums whereas Afifurrahman uses clever counting arguments and earlier results on counting integer points on modular hyperbolas.
Remark 1.3.
What about the case $r=0?$ The best known result in this direction is due to Ayyad, Cochrane, and Zhang [ACZ96] who have shown that
$\sum_{\begin{subarray}{c}x_{1}x_{2}-x_{3}x_{4}=0\\ 1\leq x_{1},x_{2},x_{3},x_{4}\leq X\end{subarray}}1=\frac{12}{{\pi}^{2}}X^{2}\log X+O\left(X^{19/{13}}(\log X)^{\frac{7}{13}}\right)$
(see [ACZ96, Thm. 3]). The main ingredient in the proof is the clever observation that
$\sum_{\begin{subarray}{c}x_{1}x_{2}-x_{3}x_{4}=0\\ 1\leq x_{i}\leq X\end{subarray}}1=\sum_{\begin{subarray}{c}(a,b)=1\\ 1\leq a,b\leq X\end{subarray}}\left(\sum_{x_{1}/x_{3}=a/b}1\right)^{2}.$
It is an interesting and challenging problem to improve on this error term.
The only other results we could find in the literature that deal with unweighted counting for indefnite forms are [GN10] and [Roe05] and both consider determinant forms. Gorodnik and Nevo (see [GN10, Cor. 2.3]) studied the problem of counting matrices in $SL(n,{\mathbb{Z}})$ inside a ball with respect to a general norm using ergodic method and in the case $n=2$, the saving in the exponent in the error term is $1/16$. Roettger [Roe05] obtained an asymptotic formula for the number of unimodular matrices over the ring of integers in a number field where the elements have height up to some parameter $t$. When specialized to our case $K={\mathbb{Q}}$, the error term gives a saving of $1/15$ in the exponent. In comparison, we save $1/2$ in the exponent of the error term in Theorem 1.1.
1.1. Some applications and further remarks
Although we are mainly motivated by the intrinsic interest in the problem, the problem of counting solutions to the determinant equations adbc=rad-bc=r, $r\neq 0$ with certain weights $\alpha(a),\alpha(b)$ etc. come up naturally, for example, in the study of amplified moments of $L$-functions (see, e.g., [DFI97], [DFI02, §16], [BC18]). In [BC18] and [DFI97], sums of the form
@@ 1.1. Some applications and further remarks
were studied and asymptotic formulae for such sums were obtained. Here, $\mathcal{N}_{i}=[N_{i}/2,N_{i}]$, $\alpha,\beta$ are arbitrary sequences, and $f,g$ are smooth functions satisfying suitable decay conditions on their derivatives. It is possible to allow arbitrary weights in some of the variables at the cost of a worse error term. For example, the following is a consequence of an intermediate step in the proof of Theorem 1.1.
@@
Our question is also related to the problem of the estimation or obtaining asymptotic formula for sums of the type $\sum_{n\leq x}\lambda(n)\lambda(n+h)$ or the corresponding smoothed sum, where $\lambda$ is an interesting arithmetical function, e.g., the divisor function, coefficients of a cusp form etc., and the shift $h$ is a fixed non-zero integer. Such sums are known in the literature as shifted convolution sums and they play a fundamental role in modern Analytic Number Theory, especially in studying moments of $L$-functions. See [Ing27], [Est32], [DI82], [Mot94], [Meu01], [DFI94], [Blo04], [HM06], to name a few. Lattice point counting problems can sometimes be interpreted as the shifted convolution problem for coefficients of appropriate automorphic forms. For example, the formula (1.1) is equivalent to the formula
@@
Estimating the error term for this sum is harder than in the case of the classical additive divisor problem. Writing $n=bc$ and $n+r=ad$, here all the four variables $a,b,c,d$ are individually restricted by sharp edges, whereas in the sum $\sum_{n\leq X}\tau(n)\tau(n+h)$, the four variables are subject to only one such restriction ($bc\leq X$) and it is possible to treat two of the three remaining variables as varying smoothly.
1.2. Discussions on the proof
Unlike in the case of counting integer points in an Euclidean ball. the choice of the norm here does not allow us to launch Harmonic Analysis directly into the counting problem. So we “forget” the group structure of $SL(2,{\mathbb{Z}})$ temporarily and treat the equation $det(\gamma)=r$ as the equation defining a surface: xyzw=rxy-zw=r. Of course, the group structure remains present throughout, only hidden, and after several transformations, it resurfaces as we are led to studying the automomorphic spectrum of $SL(2,{\mathbb{Z}})$ as explained below.
Let us first see what happens if we count with smooth weights, i.e., the case of (1.1). The first step is to eliminate one of the four variables, say $d$, and to interpret the determinant equation as a congruence modulo one of the other variables, say, $a$. Then we apply Poisson summation to one of the remaining two variables, say $b$. As usual, the zero-th frequency in the dual sum gives the main term and after one more application of Poisson summation to the remaining sum we obtain a sum of Ramanujan sums (coming from the 0-th term in new dual variable) that can be easily estimated and the following double sum: